Find the exponential function f(x) = Cat whose graph is given. f(x) = (3, 108) (1, 12)
Answer #1

we have

f(x)=Ca^x

this graph is passing through the point (1,12),

12=Ca^1

12 = Ca ............1)

and his graph is passing through the point (3,108),

108=Ca^3

108=Ca(a^2)

put Ca = 12,

108=12(a^2)

a^2=\frac{108}{12}

a^2=9

a=3

put a= 3 in equation 1),

12=C(3)

C=4

put C= 4 and a = 3 in f(x) = Cax,

f(x)=4(3^x)